### v2

# Concentration of Measure for Distributions Generated via Diffusion Models

Abstract
We show via a combination of mathematical arguments and empirical evidence that data distributions sampled from diffusion models satisfy a Concentration of Measure Property saying that any Lipschitz $1$-dimensional projection of a random vector is not too far from its mean with high probability. This implies that such models are quite restrictive and gives an explanation for a fact previously observed in the literature that conventional diffusion models cannot capture "heavy-tailed" data (i.e. data $\mathbf{x}$ for which the norm $\|\mathbf{x}\|_{2}$ does not possess a sub-Gaussian tail) well. We then proceed to train a generalized linear model using stochastic gradient descent (SGD) on the diffusion-generated data for a multiclass classification task and observe empirically that a Gaussian universality result holds for the test error. In other words, the test error depends only on the first and second order statistics of the diffusion-generated data in the linear setting. Results of such forms are desirable because they allow one to assume the data itself is Gaussian for analyzing performance of the trained classifier. Finally, we note that current approaches to proving universality do not apply to this case as the covariance matrices of the data tend to have vanishing minimum singular values for the diffusion-generated data, while the current proofs assume that this is not the case (see Subsection 3.4 for more details). This leaves extending previous mathematical universality results as an intriguing open question.
1 Introduction
One of the most astonishing contributions of deep learning is the advent of generative models for image and video generation. Diffusion-based generative models Sohl-Dickstein et al. [2015], Song and Ermon [2019],Ho et al. [2020],Song et al. [2020],Dhariwal and Nichol [2021],Song et al. [2021],Kingma et al. [2021], Karras et al. [2022], in particular, have enjoyed tremendous success in vision [LDM Rombach et al. [2022], audio [Diffwave Kong et al. [2020]] and text generation [D3PM Austin et al. [2021]]. For an overview of diffusion models and their applications, we refer to the surveys Croitoru et al. [2023] and Yang et al. [2023].
Despite significant progress in training methods, network architecture design, and hyperparameter tuning, there has been relatively little work done on understanding rigorous mathematical properties of the data generated by diffusion models. Through theory and experiments, we argue that images generated by conventional diffusion models are mathematically tractable. In fact, we argue that when the reverse process is a contraction, one can establish a concentration of measure phenomenon for the distribution of the output. Since the latter suggests a form of Gaussian Universality may hold, we compare the generalization error of linear models trained on diffusion data to the generalization error of linear models trained on Gaussian Mixtures with matching means and covariances and observe a close match.
While there are many aspects to building a diffusion model for data synthesis such as training the denoiser and choosing the forward process and noise schedule, in this work we take a higher-level approach and mainly focus on the sampling process of a pre-trained diffusion model. Our arguments are agnostic to the training procedure and the denoiser’s architectural details.
We believe that the present study is important both for advancing our theoretical understanding of generative models for images and their limitations, as well as the role of data in supervised ML:
  • We show that distributions that can be generated via diffusion models are far from arbitrary and share many properties with Gaussian distributions in a precise mathematical sense. In particular, this implies the empirical observation made in Pandey et al. [2024] that traditional diffusion models are ineffective for generating distributions with heavy tails, but this could be remedied by training the denoiser on heavy-tailed data and initiating the sampling process with heavy-tailed noise.
  • As discussed in Goldblum et al. [2023] and Nakkiran [2021], one of the major factors that hinders us from having a solid theory of deep learning is the lack of practical assumptions amenable to clean mathematical formulations and analyses regarding the true distributions of data. We believe that our work sheds light on this question in the context of image classification tasks. To elaborate further on this point, note that most image data sets encountered in practice can be approximated well using data sampled from diffusion models. This suggests that in a sense it is sufficient to explain generalization and performance of models for data coming from GMMs. The latter has been a topic of active research recently; see, e.g. Thrampoulidis et al. [2020], Loureiro et al. [2021b].
2 Related Works
The theoretical analysis of diffusion models and the images generated by such models remains an underexplored area.
  • In an emerging line of work, many papers have analyzed the output distributions and convergence of diffusion models through the lens of Langevin dynamics. Chen et al. [2022] show that with denoising diffusion probabilistic models (DDPM) and critically damped Langevin Dynamics (CLD), one can efficiently sample from any arbitrary distribution, assuming accurate score estimates - an assumption central to many works in this area. In fact, this work was among the first to provide quantitative polynomial bounds on convergence. Unfortunately, given the high-dimensional nature of the problem, estimating the score function may be practically impossible. Furthermore, it is infeasible to verify the validity of this assumption, as we do not have access to the true score function. And, as evident from the bounds of Mousavi-Hosseini et al. [2023], generating heavy-tailed distributions using Langevin dynamics initialized from the Gaussian distribution is intractable in practice as one needs to run the Langevin dynamics for an exponential number of steps. We refer to Li et al. [2024a] for a brief overview of the existing works on the convergence theory of diffusion models.
  • Seddik et al. [2020] show a form of equivalence between representations generated from Generative Adversarial Networks (GANs) and from GMMs. They considered the Gram matrix of pre-trained classifier representations of the GAN-generated images and show that asymptotically, it possesses the same distribution of eigenvalues as the Gram matrix of Gaussian samples with matching first and second moments.
  • Loureiro et al. [2021b] investigated the generalization error of linear models for binary classification with logistic loss and $\ell_{2}$ regularization. On MNIST and Fashion MNIST, they observed a close match between the real images and the corresponding GMM for the linear models and in the feature map of a two-layer network. Loureiro et al. [2021a] considered a student-teacher model and verified universality for the aforementioned datasets via kernel ridge regression. They also explored the output of a deep convolutional GAN (dcGAN), labeling it using a three-layer teacher network. Using logistic regression for classification illustrated a close match with GMMs on GAN-generated data, but also a deviation from real CIFAR10 images. Goldt et al. [2022] conducted a similar set of experiments and analyzed the generalization error of the Random Features logistic regression using the Gaussian Equivalence property. Furthermore, they considered a dcGAN trained on CIFAR100 and corroborated their theoretical findings on the dataset of images generated by the dcGAN. Then Pesce et al. [2023] studied the student-teacher model for classification and empirically demonstrated the universality of the double descent phenomenon for MNIST and Fashion MNIST. They preprocessed these datasets using a random feature map, with labels generated by a random teacher, for the ridge regression and logistic classification tasks. However, they also observed that the universality of the test error fails to hold while using CIFAR10 without preprocessing with either random feature maps, wavelet scattering, or Hadamard orthogonal projection. Moreover, Dandi et al. [2024] observed that the data distributions generated by conditional GANs trained on Fashion MNIST exhibit Gaussian universality of the test error for generalized linear models. And Gerace et al. [2024] considered mixture distributions with random labels and demonstrated universality of test error of the generalized linear models. The universality part of our work can be considered as an exploration of the same phenomenon for conditional diffusion models trained on significantly larger image datasets.
  • Refinetti et al. [2023] show that SGD learns higher moments of the data as the training continues which exhibited nonuniversality of the test error with respect to the input distribution. Exploring the limitations of current universality results and conditions under which they break remains an interesting direction of research.
  • Jacot et al. [2020] and Bordelon et al. [2020] considered kernel methods for regression and corroborated their findings through experiments on MNIST and Higgs datasets, providing evidence of Gaussian universality.
  • Pandey et al. [2024] explore heavy-tailed data generation using diffusion models. They also observe that this is not possible if one passes Gaussian noise to the models as is usually done and suggest using a Student t-distribution instead. They then demonstrate numerically that their scheme works well for generating the HRRR dataset Dowell et al. [2022].
  • Li et al. [2024b] explore a connection between diffusion models and GMMs from a different point of view. They observe that if the denoisers are over-parametrized, the diffusion models arrive at the GMMs with the means and covariances matching those of the training dataset, but learn to diverge at later stages of training. Our observations imply that even though the distribution of the diffusion-generated images stops being the same as the corresponding GMM after sufficiently many training steps, they still have many properties in common.
  • Concurrent to the submission of this work, Tam and Dunson [2025] published a preprint that establishes a similar Concentration of Measure Property. While their results are valid for any generative model consisting of Lischitz operations, they mainly explore concentration properties for GANs. Our paper conducts extensive numerical experiments with diffusion and dives into the question of bounding the Lipschitz constant of the diffusion process after $N$ steps, which is crucial to ensure that the constants in the concentration inequality can be taken to be independent of $N$. Finally, the second part of Tam and Dunson [2025] considers more abstract settings, such as generative models taking general subgaussian noise as input, while in the second part of our work we proceed to study Gaussian universality for diffusion-generated data.
3 Preliminaries and the theory
We start by defining a key notion needed for our results and then move on to provide an overview of the sampling process of diffusion models. We also prove a universality result for linear models in multiclass classification tasks. We conclude this section by stating our main findings.
3.1 Concentration of Measure Property
We use the following definition of concentration. Informally, it means that the tails of the distribution decay exponentially fast. Note that it corresponds to the Lipschitz Concentration Property for $q=2$ from Seddik et al. [2020].
Definition 1 (CoM).
Given a probability distribution $x\sim\mathbb{P}$ where $\mathbf{x}\in\mathbb{R}^{d}$, we say that $\mathbf{x}$ satisfies the Concentration of Measure Property (CoM) if there exist $C,\sigma>0$ such that for any $L$-Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$ it holds that
$\mathbb{P}\left(|f(\mathbf{x})-\mathbb{E}f(\mathbf{x})|>t\right)\leq Ce^{-(\frac{t}{L\sigma})^{2}}$ (1)
The distributions satisfying CoM arise naturally in many applications and are quite ubiquitous. We appeal to the following proposition proven in Rudelson and Vershynin [2013] :
Proposition 1.
The distribution $\mathbf{x}\sim\mathcal{N}(0,\bm{\Sigma})$ satisfies the CoM property 1. Moreover, the corresponding $C=2$ and $\sigma=\|\bm{\Sigma}^{\frac{1}{2}}\|_{op}$.
If $\bm{\Sigma}=\frac{\mathbf{I}_{d}}{d}$ and $f(\mathbf{x})=\|\mathbf{x}\|_{2}$, then Proposition 1 implies the classical fact that the norm of the normalized standard vector converges to $1$ in probability as $d\to\infty$ because in this case the upper bound of Definition 1 becomes $2e^{-(\frac{t}{\sigma})^{2}}=2e^{-td}\to 0$. However, if $\bm{\Sigma}$ is also normalized as $\text{Tr}(\bm{\Sigma})=1$ but $\|\bm{\Sigma}\|_{op}=\Theta(1)$ (which happens, for instance, if the ordered eigenvalues of $\bm{\Sigma}$ follow the power law $\lambda_{i}=Ci^{-\alpha}$ for some $C>0$ and $\alpha>1$), then the variance of $\mathbf{x}$ does not have to go to zero anymore, but Definition 1 still implies that $\mathbf{x}$ has exponentially decaying tails (to be more precise, $\mathbf{x}$ is a sub-Gaussian random vector–see Definition 3.4.1 in Vershynin [2018]) and in particular cannot be heavy-tailed. The latter was empirically demonstrated via an extensive set of experiments in Pandey et al. [2024] and therefore our results can be considered to be a theoretical validation of the corresponding body of simulations presented in Pandey et al. [2024]. Gaussians are far from the only distributions satisfying CoM; other examples include the strongly log-concave distributions, and the Haar measure-we refer to Section 5.2 in Vershynin [2018] for more examples. The concentration of measure phenomenon has played a key role in the development of many areas such as random functional analysis, compressed sensing, and information theory.
3.2 Diffusion
We provide an overview of diffusion models pertinent to our results in this paper. Given samples $x_{0}\sim q_{0}$ from a high-dimensional distribution in $\mathbb{R}^{d}$, we learn a distribution $p_{\theta}\approx q_{0}$ that allows easy sampling. A trained diffusion model essentially applies a sequence of nonlinear mappings (specifically, denoisers, denoted by $D_{\theta}$) to a white Gaussian input to obtain clean images. Following the formulation in Karras et al. [2022], assuming the distribution of the training to be "delta dirac", the score function can be expressed in terms of the ideal denoiser that minimizes $L_{2}$ error for every noise scale, i.e. $\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma)=(D_{\theta}(\mathbf{x};\sigma)-\mathbf{x})/\sigma^{2}$. This serves as a heuristic for using $(D_{\theta}(\mathbf{x};\sigma)-\mathbf{x})/\sigma^{2}$ as a surrogate for the score function to run the backward process. In most applications, $D_{\theta}$ is a neural network trained to be a denoiser, typically using a U-Net backbone. The specific denoiser we consider for our experiments is from ADM Dhariwal and Nichol [2021] which uses a modified U-Net backbone with self-attention layers. During training, the network sees multiple noise levels, and learns to denoise the images at many scales. Our analysis and statements in Section 3.5 hold for most of the diffusion models used in practice, as they employ a Lipschitz neural network. In the view of the discussion above, and setting $\sigma(t)=t$, we can represent the sampling process as an iterative procedure of $N$ steps, expressed as:
$\begin{split}&\mathbf{x}_{0}\approx\mathbf{x}^{(N)}=\mathcal{R}^{(N-1)}_{D_{\theta}}\Biggl{(}\mathcal{R}^{(N-2)}_{D_{\theta}}\biggl{(}\\ &\Bigl{(}\ldots\mathcal{R}^{(0)}_{D_{\theta}}\bigl{(}\mathbf{x}^{(0)},t_{0:1}\bigr{)}\ldots\Bigr{)},t_{N-2:N-1}\biggr{)},t_{N-1:N}\Biggr{)}\end{split}$ (2)
Where $\mathbf{x}_{T}:=\mathbf{x}^{(0)}\sim\mathcal{N}(0,t_{0}^{2}\mathbf{I})$ is isotropic Gaussian noise and $\mathbf{x}_{0}\approx\mathbf{x}^{(N)}$ is the clean image. We adopt this notation of sub-scripting the time index for $\mathbf{x}$ while super-scripting its sampler step index in order to avoid confusion with the standard notation in diffusion model papers. At any sampler step $i$ we have a (noisy image, noise level) pair $(\mathbf{x}^{(i)},t_{i})$ and the next noise level $t_{i+1}$; and $\mathcal{R}_{D_{\theta}}^{(i)}$ represents the mapping used to generate a less noisy sample i.e. $\mathbf{x}^{(i+1)}\leftarrow\mathcal{R}_{D_{\theta}}^{(i)}(\mathbf{x}^{(i)},t_{i:i+1})$, which takes in an independent noise $\bm{\epsilon}_{i}$ at each time step, as illustrated by Figure 1.
Figure 1: High-level overview of the sampling process in the diffusion models
Define $f(\mathbf{x},t):=\left(\mathbf{x}-D_{\theta}(\mathbf{x},t)\right)/t$ ;
  $\bullet$ Prob Flow ODE
1 Sample $\mathbf{x}^{(0)}\sim\mathcal{N}(0,t_{0}^{2}\mathbf{I})$
2 for  $i\in\{0,\ldots,N-1\}$  do
3       Sample $\bm{\epsilon}\sim\mathcal{N}(0,S_{\text{noise}}^{2}\mathbf{I})$
       $\hat{\mathbf{x}}^{(i)}\leftarrow\mathbf{x}^{(i)}+t_{i}\sqrt{\gamma(2+\gamma)}\bm{\epsilon}$ ;
        $\bullet$ Inject Noise
       $h\leftarrow t_{i+1}-t_{i}(1+\gamma)$ ;
        $\bullet$ Step size
       $\mathbf{x}^{(i+1)}\leftarrow\hat{\mathbf{x}}^{(i)}+hf(\hat{\mathbf{x}}^{(i)},t_{i}(1+\gamma))$ ;
        $\bullet$ Euler step
4       if $t_{i+1}\neq 0$ then
             $\mathbf{x}^{(i+1)}\leftarrow\hat{\mathbf{x}}^{(i)}+\frac{h}{2}\left(f(\hat{\mathbf{x}}^{(i)},t_{i}(1+\gamma))+f(\mathbf{x}^{(i+1)},t_{i+1})\right)$ ;
              $\bullet$ Second-order correction
5
6
return $\mathbf{x}^{(N)}$
Algorithm 1 EDM Stochastic Sampler Karras et al. [2022]
In this work, we focus on the framework considered in Karras et al. [2022] and we observe that in summary, $\mathbf{x}^{(i+1)}\leftarrow\mathcal{R}_{D_{\theta}}^{(i)}(\mathbf{x}^{(i)},t_{i:i+1})$, with
$\mathcal{R}_{D_{\theta}}^{(i)}(\mathbf{x}^{(i)},t_{i:i+1}):=\hat{\mathbf{x}}^{(i)}$
$+\frac{h}{2t_{i+1}}\left[\hat{\mathbf{x}}^{(i)}+(h+t_{i+1})d_{i}-\underbrace{D_{\theta}(\hat{\mathbf{x}}^{(i)}+hd_{i},t_{i+1})}_{\text{Denoiserafter Euler step}}\right]$
Where $d_{i}:=f(\hat{\mathbf{x}}^{(i)},t_{i}(1+\gamma))$ is as defined in line 1 of Algorithm 1 and $\gamma$ is a hyperparameter controlling the amount of additional injected noise whose scale is determined by the $S_{\text{noise}}$ hyperparameter. And $\hat{\mathbf{x}}^{(i)}$ is the current image with the added noise. Formally, we would like to claim that the distribution of the output $\mathbf{x}^{(N)}$ satisfies CoM, and we visualize the evolution of the norms of these quantities through the sampling process in Figure 4 to further illuminate our argument about the $1$-Lipschitznes of the generative process.
3.3 Classification and Gaussian Universality
We cover Gaussian universality in the context of linear multiclass classification following the framework described in Ghane et al. [2024] and extend it to an arbitrary number of classes. As we will see, most known Gaussian universality results operate in an idealized setting that does not appear to be applicable to the covariance matrices estimated from the diffusion-generated images (Figure 9). Nevertheless, we observe empirically that universality holds in the latter setting as well, hence raising a challenge of relaxing the assumptions of the existing universality results to make them more practical. We outline the corresponding notation and challenge below.
  • Consider data $\mathbf{x}\in\mathbb{R}^{d}$ being generated according to a mixture distribution with $k$ classes $\mathbb{P}=\sum_{i=1}^{k}\theta_{i}\mathbb{P}_{i}$ for $0\leq\theta_{i}\leq 1$ and $\sum_{i=1}^{k}\theta_{i}=1$. For a sample $\mathbf{x}$ from $\mathbb{P}_{i}$, i.e the i’th class, we assign a label $\mathbf{y}\in\mathbb{R}^{k}$, to be $\mathbf{y}:=\mathbf{e}_{i}$ (one-hot encoding).
    We consider a linear classifier $\mathbf{W}\in\mathbb{R}^{d\times k}$ with columns $\mathbf{w}_{\ell}$ for $\ell\in[k]$ , where for a given datapoint $\mathbf{x}$, we classify $\mathbf{x}$ based on
    $\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{w}_{\ell}^{T}\mathbf{x}$
    The generalization error of a classifier $\mathbf{W}$ on this task is defined as follows:
    $\sum_{i=1}^{k}\theta_{i}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{w}_{\ell}^{T}\mathbf{x}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}$
  • Given a training dataset $\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{n}$ with $n$ samples, where each class has $n_{i}\approx\theta_{i}n$ samples, we construct the data matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ and label matrix $\mathbf{Y}\in\mathbb{R}^{n\times k}$
    $\mathbf{X}=\begin{pmatrix}\mathbf{x}_{1}^{T}\\ \mathbf{x}_{2}^{T}\\ \vdots\\ \mathbf{x}_{n}^{T}\end{pmatrix},\quad\mathbf{Y}=\begin{pmatrix}\mathbf{y}_{1}^{T}\\ \mathbf{y}_{2}^{T}\\ \vdots\\ \mathbf{y}_{n}^{T}\end{pmatrix}$
    Without loss of generality, we can rearrange the rows of $\mathbf{X}$ to group samples from the same class together.
    We also consider a Gaussian matrix $\mathbf{G}\in\mathbb{R}^{n\times d}$ whose rows have the same mean and covariances of the corresponding rows in $\mathbf{X}$. We sometimes refer to this statement as $\mathbf{G}$ matching $\mathbf{X}$. In other words, $\mathbf{G}$ is a matrix of data sampled from the Gaussian mixture model (GMM) defined via $\sum_{i=1}^{k}\theta_{i}\mathcal{N}(\bm{\mu}_{i},\bm{\Sigma}_{i})$ where $\bm{\mu}_{i}=\mathbb{E}_{\mathbb{P}_{i}}\mathbf{x}$ and $\bm{\Sigma}_{i}=\mathbb{E}_{\mathbb{P}_{i}}\mathbf{x}\mathbf{x}^{T}-\bm{\mu}_{i}\bm{\mu}_{i}^{T}$ for $\mathbf{x}$ belonging to class $i$.
    To train for $\mathbf{W}$, we minimize $\|\mathbf{Y}-\mathbf{X}\mathbf{W}\|_{F}^{2}$ by running SGD with a constant stepsize. By the implicit bias property of SGD Gunasekar et al. [2018], Azizan and Hassibi [2018] for linear models, we observe that the iterations of SGD initialized from some $\mathbf{W}_{0}$ converge to the optimal solution of the following optimization problem
    $\min_{\mathbf{W}\in\mathbb{R}^{d\times k}}\|\mathbf{W}-\mathbf{W}_{0}\|_{F}^{2}$ (3)
    $s.t\quad\mathbf{X}\mathbf{W}=\mathbf{Y}$ (4)
    Then it is known that under the list of technical Assumptions 1 listed below the $\mathbf{W}$ obtained through running SGD on the data matrix $\mathbf{X}$ has asymptotically the same performance (generalization error) as a $\tilde{\mathbf{W}}$ obtained through running SGD on the corresponding Gaussian matrix $\mathbf{G}$, that is $\tilde{\mathbf{W}}$ solving the following optimization problem:
    $\min_{\tilde{\mathbf{W}}\in\mathbb{R}^{d\times k}}\|\tilde{\mathbf{W}}-\mathbf{W}_{0}\|_{F}^{2}$
    $s.t\quad\mathbf{G}\tilde{\mathbf{W}}=\mathbf{Y}$
In other words,
Theorem 1.
The following holds asymptotically under Assumptions 1 :
$\Biggl{|}\sum_{i=1}^{k}\theta_{i}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell}^{T}\mathbf{x}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}$
$-\sum_{i=1}^{k}\theta_{i}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\tilde{\mathbf{W}}_{\ell}^{T}\mathbf{g}\Bigl{|}\mathbf{g}\sim\mathcal{N}(\bm{\mu}_{i},\bm{\Sigma}_{i}))\Biggr{|}\to 0$
Proof.
See Appendix A. ∎
The required assumptions are as follows:
Assumptions 1.
Let $\mathbf{x}$ be any row of $\mathbf{X}$ and $\bm{\mu}$ be its mean. Then:
  • $\|\bm{\mu}\|_{2}=O(1)$
  • For any deterministic vector $\mathbf{v}\in\mathbb{R}^{d}$, and $q\in\mathbb{N}$, $q\leq 6$, there exists a constant $K>0$ such that $\mathbb{E}_{\mathbf{x}}|(\mathbf{x}-\bm{\mu})^{T}\mathbf{v}|^{q}\leq K\frac{\|\mathbf{v}\|_{2}^{q}}{d^{q/2}}$
  • For any deterministic matrix $\mathbf{C}\in\mathbb{R}^{d\times d}$ of bounded operator norm we have $Var(\mathbf{x}^{T}\mathbf{C}\mathbf{x})\rightarrow 0$ as $d\rightarrow\infty$
  • $s_{\min}(\mathbf{X}\mathbf{X}^{T})=\Omega(1)$ with high probability where $s_{\min}(.)$ is the smallest singular value.
3.4 Limitations of current universality results
Assumptions 1 hold, for example, for any sub-Gaussian $\mathbf{x}$ with mean and covariance satisfying $\|\bm{\mu}\|_{2}=O(1)$ and $\frac{c\mathbf{I}_{d}}{d}\leq\bm{\Sigma}^{\frac{1}{2}}\leq\frac{C\mathbf{I}_{d}}{d}$ (see Remark 5 in Ghane et al. [2024] for details). However, assuming that $\frac{c\mathbf{I}_{d}}{d}\leq\bm{\Sigma}^{\frac{1}{2}}\leq\frac{C\mathbf{I}_{d}}{d}$ is crucial here, as otherwise one can take a Gaussian $\mathbf{x}$ with $\bm{\Sigma}=diag(1,\frac{1}{4},\dots,\frac{1}{d^{2}})$ and $\bm{\mu}=0$ and notice that $Var(\|\mathbf{x}\|^{2})=\text{Tr}(\bm{\Sigma}^{2})-\text{Tr}(\bm{\Sigma})^{2}$ converges to a strictly positive number, violating the third part of Assumptions 1 for $\mathbf{C}=\mathbf{I}_{d}$, while $\mathbf{x}$ is normalized correctly in the sense that $\mathbb{E}_{\mathbf{x}}\|\mathbf{x}\|^{2}=\text{Tr}(\bm{\Sigma})=O(1)$.
Unfortunately, as can be seen in Figures 9 and 10, the spectra of diffusion-generated images look qualitatively similar to the "power law" $\bm{\Sigma}=diag(1,\frac{1}{4},\dots,\frac{1}{d^{2}})$, meaning that Theorem 1 does not apply in this setting. Moreover, to the best of the authors’ knowledge, such covariance matrices break the assumptions commonly made in papers focusing on universality for regression, which is usually simpler to study. For example, Montanari and Saeed [2022] also have to assume $\frac{c\mathbf{I}_{d}}{d}\leq\bm{\Sigma}^{\frac{1}{2}}\leq\frac{C\mathbf{I}_{d}}{d}$ to get concrete results for over-parametrized regression (cf. Theorem 5 in Montanari and Saeed [2022]). Despite this, as can be seen in the next section, the universality of the classification error does not break thus posing an interesting challenge of relaxing Assumptions 1 in Theorem 1.
While, technically speaking, universality is proven only for the objectives of the form (3), in practice one usually adds a softmax function $S(\mathbf{z}_{1},\dots,\mathbf{z}_{k})=(\dots,\frac{e^{\mathbf{z}_{\ell}}}{\sum e^{\mathbf{z}_{i}}},\dots)$
$\min_{\mathbf{W}\in\mathbb{R}^{d\times k}}\|\mathbf{W}-\mathbf{W}_{0}\|_{F}^{2}$ (5)
$s.t\quad S(\mathbf{X}\mathbf{W})\approx\mathbf{Y}$
Here, the approximate equality comes from the fact that the coordinates of the range of the softmax cannot turn exactly into zeros but will be very close to it on the training data if one fits the objective (5). Since this objective is of much greater practical interest than (3) and has better convergence properties, we add softmax into the objective for numerical validation of universality in the next section. Note that, from theoretical standpoint, it raises the question of incorporating softmax into the framework of Theorem 1.
3.5 Main Results
To explain why diffusion models do not perform well at generating heavy-tailed data, we prove the following result:
Theorem 2.
Assume that the denoiser $D_{\theta}(\mathbf{x}^{(i)},t_{i:i+1})$ is trained in such a way that $\|\mathcal{R}^{(i)}_{D_{\theta}}(\mathbf{x}^{(i)},t_{i:i+1})\|_{2}\leq\|\mathbf{x}^{(i)}\|_{2}$ under the notation from (2) holds for every sampling step with high probability w.r.t the randomness in $\bm{\epsilon}_{i}$. Then the resulting output $\mathbf{x}^{(N)}$ satisfies the CoM property from Definition 1 for $C=2$ and $\sigma=t_{0}$, where $\mathbf{x}^{(0)}\sim\mathcal{N}(0,t_{0}^{2}\mathbf{I}_{d})$.
Proof.
See Appendix B. ∎
The assumption $\|\mathcal{R}_{D_{\theta}}^{(i)}(\mathbf{x},t_{i:i+1})\|_{2}\leq\|\mathbf{x}\|_{2}$ might come out as very specific. In addition, it was not clear to us how to analytically verify that it is true. Nevertheless, we justify it by the following empirical observation. Understanding mathematically why Empirical Observation 1 holds thus poses an interesting challenge as well.
Empirical Observation 1.
Each sampling step $\mathbf{x}^{(i+1)}=\mathcal{R}^{(i)}_{D_{\theta}}(\mathbf{x}^{(i)},t_{i:i+1})$ of the Algorithm 1 decreases norms, i.e. $\|\mathbf{x}^{(i)}\|\leq\|\mathbf{x}^{(i-1)}\|$ is satisfied throughout the reverse process. The results of the corresponding experiments can be found in Figure 4.
We observe this contractivity in the sampling process for the setting described in Section 3.2. This observation raises the possibility that many diffusion models used in practice may also possess a contractive sampling process. An important future direction is to understand the conditions under which this property holds, given the base and target distributions, noise schedule, and denoiser training.
As explained in Subsection 3.3, CoM property is insufficient for concluding universality from any of the known universality theorems unless the upper bound from the right hand side of Definition 1 goes to 00. Nevertheless, our experiments suggest that universality holds for diffusion-generated images despite this technicality. As such, we would like to report it as an empirical observation and present the question of extending Theorem 1 to capture more complicated covariance matrices $\bm{\Sigma}$ such as the power law as an open question for future theory works.
Empirical Observation 2.
The distributions of images generated via EDM diffusion models satisfy Gaussian Universality of the test error in the sense of the conclusion of Theorem 1 for weights trained via minimizing $\|\mathbf{Y}-S(\mathbf{X}\mathbf{W})\|_{2}^{2}$ using SGD. The experiments are presented in Figure 6 preceded by the description of the setup.
4 Experiments
We conduct a series of experiments where we train linear classifiers on diffusion-generated images and on samples from a Gaussian Mixture Model (GMM) with matching means and covariances. We also empirically investigate the concentration properties of these images. Throughout all our experiments, we use the trained conditional diffusion model checkpoint from EDM Karras et al. [2022], which uses the ADM architecture Dhariwal and Nichol [2021] and was trained on Imagenet64 (Imagenet-1k Deng et al. [2009] downscaled to $64\times 64$ pixels). Moreover, we sample according to the EDM stochastic sampler outlined in 1, using the settings recommended by the authors.
[…]
Figure 2 presents some samples from our dataset. See Appendix Figure 20 for more samples and Figure 9 for the spectra of the covariance matrices per class.
Figure 2: Samples from conditional diffusion model with ADM architecture Dhariwal and Nichol [2021] using the checkpoint and sampler settings from Karras et al. [2022], trained on Imagenet64.
Figure 3: The distribution of the $\ell_{2},\ell_{4}$ and $\ell_{10}$ norms of diffusion-generated images for 20 classes of Imagenet64 in our experiments, computed over 10240 samples per class.
4.1 Norm Evolution
We empirically investigate the concentration of norms throughout the sampling process. Following the recommendations of Karras et al. [2022], the diffusion sampling process of $N=256$ steps, with $\sigma(t)=t$, begins with isotropic Gaussian noise of scale $t_{max}:=t_{0}=80$. The noise schedule is constructed as $t_{i<N}=\left({t_{max}}^{\frac{1}{\rho}}+\frac{i}{N-1}\left({t_{min}}^{\frac{1}{\rho}}-{t_{max}}^{\frac{1}{\rho}}\right)\right)^{\rho}$ with $t_{min}:=t_{N-1}=0.002$ and final noise scale $t_{N}=0$. Here, $\rho=7$ is a hyperparameter observed to improve image quality.
Figure 4: The evolution of the $\ell_{2}$ norms through the stochastic sampling process 1. The noise schedule and sampler settings are as prescribed in Karras et al. [2022].
Figure 5: The difference in $\ell_{2}$ norms of intermediate images between consecutive steps of the EDM sampling process. Figure (a) shows the evolution of $||\mathbf{x}^{(i-1)}||-||\mathbf{x}^{(i)}||$ ($i$ denotes sampling step) vs. noise scale for 5000 generation trajectories across 5 classes and (b) shows the mean with variance envelopes of the trajectories, computed over 2048 samples per class. The sampling process clearly appears to be a contraction, supporting Empirical Observation 1
We present further empirical observations regarding the sampling process in Appendix Section D. In Figure 13 we show how the sampling process progressively matches the eigenvalues of the Gram matrix. In Figure 14 we investigate the evolution of the norms of individual pixels.
4.2 Linear Classifier Experiments
We train linear classifiers on our dataset of diffusion-generated images and on the corresponding Gaussian data sampled from a GMM fitted on 10240 diffusion-generated images per class. Following the setting of subsection 3.3, we use SGD as our optimizer and mean squared error (MSE) as our loss criterion. For multi-class classification, we use a softmax activation on the logits and compute the MSE loss against the one-hot-encoded class labels. For binary classification, we compute the MSE loss on the logit after sigmoid activation. This regression on predicted class probabilities is done in practice when working with soft (noisy) labels or in the context of knowledge distillation.
We compare the accuracies achieved by linear classification on the diffusion-generated images versus the GMM samples, when varying the size of the training set between 128 and 4096 samples per class. Figures 6 and 7 present the results of 10-20 independent runs per training data split, with a different random seed for each run. Thus, for each run, a unique pseudorandom generator state determines weight initialization in addition to the sampling and minibatch shuffling of $N_{\text{train per class}}\in[128,4096]$ samples from our dataset of 10240 samples per class for diffusion-generated images and GMM samples. Likewise, we fix the size of our held-out test set to $N_{\text{test per class}}=1024$, randomly sampled according to each run’s unique random state from a separate subset of our dataset to ensure no overlap with the training set.
Figure 6: Linear classifier accuracies for diffusion-generated images (Red) and GMM samples (Blue). We consider 4, 10, and 20 class mixtures of Baseball, Cauliflower, Church, Coral Reef, English Springer, French Horn, Garbage Truck, Goldfinch, Kimono, Mountain Bike, Patas Monkey, Pizza, Planetarium, Polaroid, Racer, Salamandra, Tabby, Tench, Trimaran, Volcano.
Figure 7: Binary linear classifier accuracies for diffusion-generated images (Red) and GMM samples (Blue).
To robustly achieve the best possible linear classification performance, we perform an extensive sweep over a range of batch sizes and of learning rates between $[10^{-4},0.1]$, while ensuring convergence with respect to the test loss, and no overfitting. For practical reasons, we use a cosine annealing learning rate schedule to speed up convergence.
In summary, we observe matching accuracies for linear classifiers trained on diffusion-generated images and on the corresponding Gaussian data, for a range of training set sizes and class subsets. At large values of $N_{\text{train per class}}$ we begin to see a divergence in the accuracies, which we attribute to the estimation gap from computing the mean and covariance from a finite number of samples. The choice of MSE loss on one-hot-encoded labels may seem unconventional for classification but is done to match the setting of 3.3. We re-ran our experiments using cross-entropy loss and also observe a match, but did not conduct as extensive a sweep for this setting and expect the match to improve.
Figure 8: Accuracies for linear classifier trained with cross-entropy loss, for diffusion images (Red) and GMM (Blue).
In Appendix Section C we compute the eigenvalue spectra of the Gram matrices of multi-class mixtures of diffusion-generated images. Figures 11 and 12 in the appendix show a very close match between the Gram spectra of diffusion-generated images and that of the corresponding GMM. And in Appendix Section E we investigate the Gram spectra and eigenspaces of ResNet representations of high-resolution images sampled from a latent-diffusion model, showing a close match (Figures 15, 16, and 17) and suggesting an investigation into the concentration of latent diffusion models and their representations as an avenue for future work.
5 Conclusion
In this work, we focus on characterizing the mathematical properties of images generated by diffusion models by examining the generalization error of the classification tasks. We are motivated by the fact that characterizing the generalization error and performance of neural networks precisely remains one of most challenging problems in modern machine learning. In fact, most theoretical works have focused on analyzing models under specific assumptions about data distribution, such as isotropic Gaussianity even though real-world datasets are almost never Gaussian. As such, we choose to study theoretical properties of the diffusion-generated distributions instead as an approximation to real-world distributions more amenable to analyses. Future directions include extending the universality results to accommodate for more general covariance matrices, incorporating training with softmax into the universality framework and providing a rigorous proof of the contractivity of the sampling process.
Figure 9: Spectra of covariance matrices of diffusion-generated images. Scaled by exponent $0.1$ for presentation.
Figure 10: First 1000 eigenvalues of the covariance matrices.
6 Acknowledgments
We would like to thank Joel A. Tropp for insightful discussions and Morteza Mardani for suggesting to look into the evolution of the norms of the individual pixels during the sampling process. AB is grateful for the support of the Basdall Gardner Memorial MCD Fellowship and for the gracious maintenance of computational resources carried out by the Research Computing Task Force (RCTF) at UT Austin.
Appendix A Proof of Theorem 1
First note that we will utilize a multi-dimensional version of the CLT result of Bobkov [2003] (Corollary 2.5) which controls the following quantity for a matrix $\mathbf{W}$ with "generic" column vectors:
$\biggl{|}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{X})}^{T}\mathbf{x}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}-\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{X})}^{T}\mathbf{g}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}\biggr{|}$
This generalization follows by applying a union bound argument. For the main quanitity of interest,
$\biggl{|}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{X})}^{T}\mathbf{x}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}-\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{G})}^{T}\mathbf{g}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}\biggr{|}$
We would need to bound the following:
$\biggl{|}\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{X})}^{T}\mathbf{g}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}-\mathbb{P}\Bigl{(}i\neq\operatorname*{arg\,max}_{\ell\in[k]}\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{G})}^{T}\mathbf{g}\Bigl{|}\mathbf{x}\sim\mathbb{P}_{i}\Bigr{)}\biggr{|}$
Which involves analyzing the covariance and the mean of $\mathbf{W}_{\ell,\Phi_{\lambda}(\mathbf{A})}^{T}\mathbf{g}$ for $\mathbf{A}=\mathbf{G},\mathbf{X}$.
We know from Thrampoulidis et al. [2020] for the case of a GMM, (see Equation 2.7 in Thrampoulidis et al. [2020]) that the generalization error is characterized by the quantities $\bm{\mu}_{\ell}^{T}(\mathbf{w}_{\ell}-\mathbf{w}_{\ell^{\prime}})$, and $\bm{\Sigma}^{1/2}_{\ell}\mathbf{S}\bm{\Sigma}^{1/2}_{\ell}$ where $(S_{\ell})_{ij}:=(\mathbf{w}_{i}-\mathbf{w}_{j})^{T}(\mathbf{w}_{i}-\mathbf{w}_{j})$ for $i,j\neq\ell$. Consider the following ridge regression objective:
$\Phi_{\lambda}(\mathbf{A}):=\min_{\mathbf{W}}\frac{\lambda}{2}\|\mathbf{A}\mathbf{W}-\mathbf{Y}\|_{F}^{2}+\|\mathbf{W}\|_{F}^{2}=\sum_{\ell=1}^{k}\min_{\mathbf{w}_{\ell}}\frac{\lambda}{2}\|\mathbf{A}\mathbf{w}_{\ell}-\mathbf{y}_{\ell}\|_{2}^{2}+\|\mathbf{w}_{\ell}\|_{2}^{2}$
We denote the solution to the above optimization problem as $\mathbf{W}_{\Phi_{\lambda}(\mathbf{A})}$. Now in order to characterize $(S_{\ell})_{ij}$, note that we need to understand the pairwise interaction of $\mathbf{w}_{i}$ and $\mathbf{w}_{j}$ and the decomposition provided cannot capture these quantities. To do this, we use the following identity:
$\min_{\mathbf{w}_{i},\mathbf{w}_{j}}\frac{\lambda}{2}\|\mathbf{A}\mathbf{w}_{i}-\mathbf{y}_{\ell}\|_{2}^{2}+\|\mathbf{w}_{i}\|_{2}^{2}+\frac{\lambda}{2}\|\mathbf{A}\mathbf{w}_{j}-\mathbf{y}_{\ell}\|_{2}^{2}+\|\mathbf{w}_{j}\|_{2}^{2}$
$=\min_{\mathbf{w}_{i}-\mathbf{w}_{j},\mathbf{w}_{i}+\mathbf{w}_{j}}\frac{\lambda}{4}\|\mathbf{A}(\mathbf{w}_{i}+\mathbf{w}_{j})-\mathbf{y}_{\ell}\|_{2}^{2}+\|\mathbf{w}_{i}+\mathbf{w}_{j}\|_{2}^{2}+\frac{\lambda}{4}\|\mathbf{A}(\mathbf{w}_{i}-\mathbf{w}_{j})-\mathbf{y}_{\ell}\|_{2}^{2}+\|\mathbf{w}_{i}-\mathbf{w}_{j}\|_{2}^{2}$
And by studying the norms of $\bm{\Sigma}_{\ell}^{1/2}(\mathbf{w}_{i}\pm\mathbf{w}_{j})$ we can understand $(S_{\ell})_{ij}$. Note that in the argument above $\mathbf{A}$ could be either $\mathbf{X}$ and $\mathbf{G}$ as a multi-dimensional CLT argument reduces the problem of universality of the test error on $\mathbf{X}$ to $\mathbf{G}$ and it only requires the first and second order statistics of $\mathbf{X}$. Note that by combining the results of Ghane et al. [2024] and the above identity, we observe that $|(S_{\ell}(\mathbf{X}))_{ij}-(S_{\ell}(\mathbf{G}))_{ij}|\xrightarrow[]{\mathbb{P}}0$ for every $\lambda$. We observe that if $\Phi_{\lambda}(\mathbf{A})\xrightarrow[]{\mathbb{P}}c_{\lambda}$ then $\sup_{\lambda>0}\Phi_{\lambda}(\mathbf{A})\xrightarrow[]{\mathbb{P}}\sup_{\lambda>0}c_{\lambda}$. Combining this with a perturbation argument concludes the proof.
Appendix B Proof of Theorem 2
We rely on the following proposition, claiming that the image of an exponentially concentrated distribution under a Lipschitz transformation remains exponentially concentrated, albeit with different constants. It follows trivially from the definition of concentration but is nonetheless crucial for our purposes:
Proposition 2.
Let $\mathbf{x}$ satisfy Definition 1 with constants $C,\sigma$ and assume that $\mathcal{G}:\mathbb{R}^{d}\to\mathbb{R}^{D}$ is $L$- Lipschitz. Then $\mathcal{G}(\mathbf{x})$ satisfies CoM from Definition 1 as well with constants $C,L\sigma$.
Now, recall that the images are generated step by step according to:
$\mathbf{x}_{0}\approx\mathbf{x}^{(N)}=\mathcal{R}^{(N-1)}_{D_{\theta}}(\mathcal{R}^{(N-2)}_{D_{\theta}}(\ldots\mathcal{R}^{(0)}_{D_{\theta}}(\mathbf{x}^{(0)},t_{0:1})\ldots),t_{N-2:N-1}),t_{N-1:N})$ (6)
We will prove that each $\mathbf{x}_{i}$ satisfies the CoM of measure property by induction by $i=0,\dots,N$:
$\bullet$ Basis $i=0$ follows from Proposition 1
$\bullet$ Step $i\to i+1$ would follow by applying Proposition 2 with $L=1$ to $\mathbf{x}^{(i+1)}=\mathcal{R}^{(i)}_{D_{\theta}}(\mathbf{x}^{(i)},t_{i:i+1})$ if we knew that $\mathcal{R}^{(i)}_{D_{\theta}}(.,t_{i:i+1})$ is $1$-Lipschitz. Since we already assume that $\mathcal{R}^{(i)}_{D_{\theta}}$ is norm-decreasing, it suffices to just prove that it is Lipschitz. The latter does not have to hold in general but in this case we make a specific assumption that $\mathcal{R}^{(i)}_{D_{\theta}}$ corresponds to a denoiser employing a U-Net architecture. This is a neural network consisting of the following blocks:
  • Fully-Connected Layers with a Lipschitz activation function $\sigma=SiLU$ and a matrix of weights $\mathbf{W}$. These are Lipschitz functions with constant $\|\sigma\|_{Lip}\|\mathbf{W}\|_{op}$.
  • Convolutional Layers with a filter $\mathbf{W}$. These are also Lipschitz functions with constant $\|\sigma\|_{Lip}\|\mathbf{W}\|_{op}$.
  • Self-Attention Layers As shown in Kim et al. [2021], these are not Lipschitz over the entire domain. However, it can be seen from the same derivations that, if we restrict the domain to points from a distribution satisfying CoM, then it is Lipschitz with high probability. Thus, for our specific scenario, these are Lipschitz as well.
  • Max Pool, Average Pool, Group Normalization, Positional Embedding, Upsampling and Downsampling Layers. All these layers are $1$-Lipschitz.
We conclude that the mapping $\mathbf{x}^{(i)}\to\mathbf{x}^{(i+1)}$ is Lipschitz but, technically speaking, the constant is unbounded without the assumption that $\mathcal{R}^{(i)}_{D_{\theta}}$ does not increase norms. Moreover, since the sampling process involves $N$ steps for a relatively big $N$, the Lipschitz constant of the mapping $\mathbf{x}_{T}\to\mathbf{x}_{0}$ might accumulate and explode unless the Lipschitz constant of each step is bounded by $1$. While we could not prove directly that the latter is the case so far, we observed it to be the case in the simulations we have conducted (cf. Figure 4). As such, we decided to assume that the training is performed in such a manner that the sampling steps $\mathbf{x}^{(i)}\to\mathbf{x}^{(i+1)}$ are all $1$-Lipschitz mappings for the scope of the present work. In addition, we can prove that the same conclusion holds if we only assume that each $\mathcal{R}_{i}$ is norm-decreasing for large enough $t>K$ for a constant $K$. This follows automatically from the following result, that we could not find in the literature:
Theorem 3.
If $(x,y)\sim\Pi$ where $\Pi$ is a joint distribution with marginals, $\pi_{1\#}\Pi=p_{1}$ and $\pi_{2\#}\Pi=p_{2}$ and $p_{1}$ and $p_{2}$ are two distributions satisfying CoM, then $(x,y)$ also satisfies CoM.
Proof.
The proof technique is adopted from Ledoux [2006]. We also extend that result to accommodate for arbitrary coupling on $(x,y)$. For every $L$-Lipschitz function $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, we have from triangle inequality
$\mathbb{P}(|f(x,y)-\mathbb{E}_{\Pi}f(x,y)|>2t)\leq\mathbb{P}(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t)+\mathbb{P}(|\mathbb{E}_{p_{1}}f(x,y)-\mathbb{E}_{\Pi}f(x,y)|>t)$ (7)
For the first term in 7, we have that for the joint distribution $\Pi$
$\mathbb{P}(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t)$ $=\mathbb{E}_{\Pi}\mathds{1}\{|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t\}=\mathbb{E}_{p_{2}}\mathbb{E}_{p_{1}|p_{2}}\mathds{1}\{|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t\}$
$=\mathbb{E}_{p_{2}}\mathbb{P}_{p_{1}|p_{2}}(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t)$
Now we observe that for every $y$, $f(x,y)$ is also $L$-Lipschitz in $x$, thus by CoM
$\mathbb{P}(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t)=\mathbb{E}_{p_{2}}\mathbb{P}_{p_{1}|p_{2}}(|f(x,y)-\mathbb{E}_{p_{1}}f(x,y)|>t)\leq\mathbb{E}_{p_{2}}Ce^{-(\frac{t}{L\sigma})^{2}}=Ce^{-(\frac{t}{L\sigma})^{2}}$
For the second term in 7, letting $g(y):=\mathbb{E}_{p_{1}}f(x,y)$, we observe that $g$ is also Lipschitz, so by CoM for $p_{2}$,
$\mathbb{P}(|\mathbb{E}_{p_{1}}f(x,y)-\mathbb{E}_{\Pi}f(x,y)|>t)\leq Ce^{-(\frac{t}{L\sigma})^{2}}$
Summarizing, we obtain that:
$\mathbb{P}(|f(x,y)-\mathbb{E}_{\Pi}f(x,y)|>2t)\leq 2Ce^{-(\frac{t}{L\sigma})^{2}}$
Appendix C Gram Spectrum
For each of the subsets of classes we considered for our multiclass linear classification experiments, we also investigated the spectrum of the gram matrix of the corresponding mixture distribution. Using an equal number of samples per class, we construct a data matrix $X\in\mathbb{R}^{n\times d}$ where $n$ is the total number of samples and $d=12288$ is the dimensionality of each sample (viewed as a vector). Figure 11 presents the eigenvalue spectrum of the resulting Gram matrix of the type $XX^{T}\in\mathbb{R}^{n\times n}$. As can be seen, we observe a very close match between the distributions of the eigenvalues of the Gram matrices for the diffusion-generated data and the corresponding GMM, but there is a slight mismatch for the smaller eigenvalues. We leave the question of finding out if there are any reasons for the latter mismatch apart from numerical inaccuracies for future work.
Figure 11: Spectra of Gram Matrices for balanced mixtures of samples from 4, 10, and 20 classes considered in the linear classification experiments. Computed for Diffusion (Red) and GMM (Blue). We use 2048 samples per class for 4 classes, and 512 samples per class for 10 and 20 classes.
Figure 12: First 1000 eigenvalues of Gram Matrices for balanced mixtures of 4, 10, and 20 classes.
Note that while establishing the closeness of eigenvalue distributions of the Gram matrices allows one to characterize the behavior of certain algorithms such as Least-Squares SVM or spectral clustering, this does not allow us to analyze more elaborate algorithms. For example, the LASSO objective $\min_{w}\|Xw-y\|+\lambda\|w\|_{1}$ for $w\in\mathbb{R}^{d},X\in\mathbb{R}^{n\times d}$ is not unitarily invariant. Hence, given $X^{\prime}\in\mathbb{R}^{n\times d}$, even knowing that the Gram matrices $(X^{\prime})^{T}X^{\prime}$ and $X^{T}X$ are exactly equal to each other does not let one conclude that $w^{\prime}$ identified via $\min_{w^{\prime}}\|X^{\prime}w^{\prime}-y\|+\lambda\|w^{\prime}\|_{1}$ yields to performance similar to the performance of $w$.
Appendix D Sampling Process
In Figures 4 and 5, we plotted the evolution of the norms. Now in Figure 13 we also observe that the sampling process first matches the higher eigenvalues of the Gram matrix, and then progressively matches the lower eigenvalues.
Figure 13: Eigenvalues of Gram matrix through EDM sampling process. Computed with 2048 samples of a single class.
D.1 Evolution of Norms of Pixels
We also investigate the norms of individual pixels through the EDM sampling process. Note that Figure 14 d is on a log-log scale, which cuts off negative values of the plotted standard deviation envelope at the low noise scales; indeed, at any step of the sampling process, there are pixels that increase in norm. But on average, the pixel norms are decreasing.
Figure 14: (a) shows distribution of the pixel norms for a single class, through the EDM sampling process. (b) shows the individual trajectories of the norms of 1000 randomly selected pixels at different noise scales of sampling. (c) shows the difference in norms between sampling steps, for 100 randomly selected pixels. (d) shows the mean and standard deviation envelope of this difference in norms. Note that this is on a log-log scale and negative values of $(\mu-\sigma)$ are cut off.
Appendix E Representations of High-Resolution Latent Diffusion Samples
Lastly, we investigate pre-trained classifier representations of high-resolution images generated from a latent diffusion model. We generated a dataset of $512\times 512$ px images with deterministic sampling from EDM2 Karras et al. [2024a] (large) using classifier-free guidance and guidance strength chosen to minimize Fréchet distance computed in the DINOv2 feature space Oquab et al. [2024]. We then resize to $256\times 256$ px and apply the standard $224\times 224$ px center crop before feeding to ResNets He et al. [2015] of various depths. This pre-processing is done to match the resolution that these ResNets were trained on. The representations are the output after global average pooling, before the final fully connected layer. They are of dimension 512 for Resnet18 and 2048 for Resnet50 and Resnet101.
Figure 15: Spectra of Gram Matrices of ResNet representations of a 4-class mixture (Church, Tench, English Springer, French Horn). Computed for representations of 1350 images per class from EDM2 (Red) and for GMM fitted on those representations (Blue).
Figure 16: Top eigenspaces of Gram matrices of ResNet representations of 4-class mixture of EDM2 images. Corner plot of eigenvector $i$ vs. $j$ (Gaussian KDE) for representations of EDM2 (Red) and for GMM fitted on representations (Blue).
As seen in Figures 15 and 16, the Gram matrices of the ResNet representations of diffusion images show a close to match to their GMM counterparts when viewing the eigenvalue spectrum. Compare with Figure 18, which presents the spectra of the covariance matrix of diffusion representations of a single class. Moreover, we observed clear separability of the classes in the first few eigenvectors. This motivated us to train a logistic regression model on the top eigenvectors of these Gram matrices, and as shown in Figure 17, the first 3-4 eigenvectors are all that are needed for near-perfect accuracy. We leave the question of how this scales with the number of classes for future work. We plot the mean and standard deviation envelope over 100 runs of logistic regression, each using a random split proportion of $0.8$ training samples from the 1350 representations per class in the mixture. The test set is fixed as a $0.2$ proportion of the representations of real images.
Figure 17: Logistic regression trained on ResNet representations of a 4-class mixture. Shown for EDM2 representations (Red), for GMM fitted on EDM2 representations (Blue), and for representations of real images (Green).
As a latent diffusion model, EDM2 Karras et al. [2024b] does diffusion in the latent space of a pre-trained variational autoencoder (VAE). We investigate the evolution of the norms of the latents through the deterministic sampling process. The results are presented in Figure 19.
Figure 18: Eigenvalues of covariance matrix of ResNet representations of EDM2 diffusion-generated images, for a single class, computed over 4096 samples.
Figure 19: Evolution of norms of latents through EDM2 deterministic sampling, for a single class. These are latents of dimension $4\times 64\times 64$ in the latent space of a pre-trained VAE. We refer to Karras et al. [2024b] for further details.
Figure 20: More samples from our diffusion-generated dataset, demonstrating the visual fidelity of the generated images. All 20 Imagenet64 classes used in our experiments are represented